OpenStudy (anonymous):

I've got a math exam on friday. I have some of the previos exams here and there are a whole bunch of questions which I cant solve. I really needs some help with them. I'll be posting them here now. Thank you very much for all who help me out, I really appreciate it :)

OpenStudy (anonymous):

Limits problems, find the limit: a) \[\lim_{x \rightarrow 5}x/(x-5)\] b) \[\lim_{x \rightarrow \infty}(\sqrt{x+1}-\sqrt{x})\]

OpenStudy (anonymous):

a) add and subtract 5 frm d numerator that gives u 1 + 1/(x-5) when u put in infinity it is 1 + 1/inf = 1

OpenStudy (anonymous):

as for b) multiply and divide by \[\sqrt{x+1}+\sqrt{x}\]

OpenStudy (anonymous):

that gives u 1 / √x+1+√x

OpenStudy (anonymous):

when u put in infinity it gives u 1/inf = 0

OpenStudy (anonymous):

\[a) \lim_{x \rightarrow 5} { x \over x-5} =\infty\] That's because when you substitute 5 in the top you get 5, and in the bottom you get 0. A number over 0 is infinity.

OpenStudy (anonymous):

Next up, find the derivative: a) \[f(x) =x^{2}\times e^{x ^{2}}\] b) \[f(x,y)=(y ^{3}x+e ^{2y})^{3}\]

OpenStudy (anonymous):

a) dy/dx= 2xe^x3 + 3x^4e^x^3

OpenStudy (anonymous):

b) him1618 already got the right answer.

OpenStudy (anonymous):

was a wrong?

OpenStudy (anonymous):

I guess :)

OpenStudy (anonymous):

By the way, in a) if the limit approaches 5 from the left, it would be -infinity.

OpenStudy (anonymous):

so ur saying it doesnt exist?

OpenStudy (anonymous):

Yep.

OpenStudy (anonymous):

yeah neway its a sort of hyperbola centred at 5

OpenStudy (anonymous):

It has a vertical asymptote at x=5.

OpenStudy (anonymous):

Question 3: Define the min/max of \[f(x)=x^{3}-12x ^{2}+21x+25\] in [-1, infinity[

OpenStudy (anonymous):

differentiate it f'(x) = 3x^2 - 24x + 21

OpenStudy (anonymous):

the roots are 7, 1

OpenStudy (anonymous):

Question 4: Define the local min/max of \[f(x,y)=x ^{3}-y ^{3}+3xy\]

OpenStudy (anonymous):

Sick how good you are :) Two more questions...

OpenStudy (anonymous):

yeah go on

OpenStudy (anonymous):

Question 5: Solve the following inequalities a) \[x^{3}+2x ^{2}+2x \le0\] b) \[\log_{2} x <2+\log_{4} x\]

OpenStudy (anonymous):

Question 6: Solve the following equations a) \[\left| x-3 \right|+\left| x+2 \right|=10\] b) \[2^{3x}=3^{2x}\]

OpenStudy (anonymous):

a) x<=0

OpenStudy (anonymous):

is it right

OpenStudy (anonymous):

I believe it's right since it's factored to x(x^2+2x+2). But the part x^2+2x+2 is always >0. So, the sign just depends on x. Hence the whole expression is <=0 when x<=0.

myininaya (myininaya):

anwar is like totatally right!

OpenStudy (anonymous):

so am i man..dont demean me..lol

OpenStudy (anonymous):

Why is the part x^2+2x+2 always >0?

OpenStudy (anonymous):

because its determinant is negative...

OpenStudy (anonymous):

Yeah. You're the one who got the answer. I was just following your foot steps :)

OpenStudy (anonymous):

it is an upward parabola where the min value is greater than 0

OpenStudy (anonymous):

Ah I understand :)

OpenStudy (anonymous):

You deserve a medal him1618.

OpenStudy (anonymous):

thnx a lot

OpenStudy (anonymous):

I gave one too :)

OpenStudy (anonymous):

You're welcome :)

OpenStudy (anonymous):

so kind

OpenStudy (anonymous):

I want to solve part b) in inequalities.

OpenStudy (anonymous):

b) x is in 0 to 16

OpenStudy (anonymous):

Please do I can't solve it :/

OpenStudy (anonymous):

thats d ans

OpenStudy (anonymous):

its damn tedious to write these eqns

OpenStudy (anonymous):

Yeah, that sucks....

OpenStudy (anonymous):

on d comp only..on paper it took a minute

myininaya (myininaya):

i think i considered all the cases for 6a

OpenStudy (anonymous):

2 is \[\log_{4}16 \]

OpenStudy (anonymous):

\[\log_2x-\log_4x<2 \implies \log_2x-{\log_2x \over \log_24}<2 \implies \log_2x-{1 \over 2}\log_2x<2\] \[{1 \over 2}\log_2 x<2 \implies \log_2{\sqrt x}<2 \implies \sqrt x<4 \implies x<16\]

OpenStudy (anonymous):

But x has to be a positive number, So 0<x<16

OpenStudy (anonymous):

right

OpenStudy (anonymous):

Nice answer myininaya helpd me a bunch :)

myininaya (myininaya):

here is 6b

OpenStudy (anonymous):

yeah 6a is bloody tedious..nice wrk

OpenStudy (anonymous):

b is right 2

OpenStudy (anonymous):

Right on the spot AnwarA I'll jot that down to my notebook :)

OpenStudy (anonymous):

Did anyone solve 6a, or I can take it?

OpenStudy (anonymous):

yes its been solved by myininaya

OpenStudy (anonymous):

I hate you man >.<

myininaya (myininaya):

woman* lol

OpenStudy (anonymous):

OMG.. I don't hate you at all :P

myininaya (myininaya):

lol

myininaya (myininaya):

you can look over my work. there might be a mistake who knows

myininaya (myininaya):

it looks good to me though

OpenStudy (anonymous):

I can hardly read it :P

myininaya (myininaya):

:(

myininaya (myininaya):

i can rewrite it

OpenStudy (anonymous):

No, It's right. I got the exact same answers.

myininaya (myininaya):

you can also graph it in your calculator if you want to verify your answers

OpenStudy (anonymous):

x=15/2 for x>=3, x=-5/2 for x<=2. No solution for 2<x<3.

myininaya (myininaya):

right!

myininaya (myininaya):

i can write neat when i'm not trying to beat you guys to answer the questions lol

myininaya (myininaya):

i wish in high school there was a math quiz bowl and i participated that would have been so fun

OpenStudy (anonymous):

lol we should divide the questions equally. I wish I had a scanner -.-

OpenStudy (anonymous):

One question, what happens here: \[\log_2x-{\log_2x \over \log_24}<2 \implies \log_2x-{1 \over 2}\log_2x<2\] ?

OpenStudy (anonymous):

log4 to base to is 2.

OpenStudy (anonymous):

to base 2*

myininaya (myininaya):

because 2^2=4

OpenStudy (anonymous):

Mm, yes of course! :)

OpenStudy (anonymous):

Do you still have any questions, or that's it?

OpenStudy (anonymous):

I think that is about it :) So quick and dirty. Thanks a million for all of you!

myininaya (myininaya):

yes all your questions were being answered so quickly by him hes such a meanie

OpenStudy (anonymous):

no probs mate

OpenStudy (anonymous):

yeah thats kind...

myininaya (myininaya):

lol i'm kidding

OpenStudy (anonymous):

dont sweat..i hate studying so much...bt i lov doing this

OpenStudy (anonymous):

I think that is about it :) So quick and dirty. Thanks a million for all of you!

OpenStudy (anonymous):

well thnx

OpenStudy (anonymous):

Good luck in your exam on Friday :)

OpenStudy (anonymous):

yeah good luck...remember me if u feel nervy..lol

OpenStudy (anonymous):

Thanks, I believe I can make it :D